OpenStudy (january123):

What is the solution to y = 2x + 3 for the replacement set x = {1.3}? Fan~Medal☺

OpenStudy (butterflydreamer):

similar to what you did in the previous question :) Substitute x = 1.3 into the equation " y = 2x+3" So basically where you see "x" in y = 2x + 3, replace the x with 1.3 y = 2 (1.3) + 3 y = ?

OpenStudy (january123):

what a second...the answer?